(17) Assume X and Y have the following joint density function: f(x,y) =-(x + y), 0<x<2,0<p<2 Calculate P(X +Ys 1). B) C) 17 24 A) E) 24 24 The circled answer is wrong, please show steps to arrive at a correct answer.
Answer #1

P(X+Y<=1)=\int_{0}^{1}\int_{0}^{1-y}f(x,y) dx dy =\int_{0}^{1} (1/8)*(x+y) dx dy =\int_{0}^{1} (1/8)*(x2/2+xy|1-y0 dy

=\int_{0}^{1}(1/8)*((1-y)2/2+y-y2) dy=(1/8)*((1-y)3/3+y2/2-y3/3)|10 =(1/8)*(-1/3+1/2-1/3)=1/24

option A is correct